Part A What type of industry is likely to use a process cost system? List examples Why is it important to establish appropriate cost pools?

PART B An electrical component manufacturer has selected direct labour hours as an application base. They plan to sell 35,000 units of copper tubing although the factory has the capacity to produce 40,000 units under normal circumstances.

Overheads are estimated as follows:

• Indirect materials and supplies $96,200

• Repairs and maintenance $24,100

• Plant service contracts $37,000

• Refurbishing cost $89,100

• Machinery depreciation $185,000

• Factory insurance $18,200

• Property taxes $4,500

• Heat, light and power $51,700

• Miscellaneous factory overhead $6,000

• indirect Labour $120,000.00

• Materials $80,000.00

• Transportation $25,000.00

• Rent $40,000.00

• Security Cost $15,000.00

All overhead costs except depreciation, property taxes and miscellaneous factory overhead are expected to increase by 10% during the year.

Depreciation should increase by 12% and a 20% increase in property taxes and miscellaneous overhead is expected.

A total of 55,600 direct labour hours was actually used to produce 35,000 units of copper tubing .

Direct labour hours is expected to increase to 60,000 hours as production volume increases.

a. Calculate the predetermined overhead rate

b. Determine the Total estimated overheads?

c. Calculate the overhead application rate if 60,000 Direct Labour hours was used in production .

Ascertain the overhead applied to the job if it took 120 direct labor hours.
Answer #1

In Part A,

Process costing is used where manufacturing of products are done in a large quantity and done by following multiple process. Many industries like coal mining, paper production, plastic production, food industry, chemical, textile industry, cement and soap industry or like these where a process is required to complete the manufacturing of a unit of product.

In part B,

A. Pre determined overhead rate is 21.568 which is calculated as follows.

Step 1: calculate Total of overheads

Sum of total overheads = 754800

Total units produced = 35000

Predetermined cost per unit = Sum of overheads\ number of units produced

= 754800/35000

= 21.568 $ per unit

=754800/ 55600

= 13.58 $ per hour

B. Calculation of total Estimated Overheads

Cost sheet for calculation of overhead.

Depreciation ( 185000+12℅) = 207200

Property tax (4500+20℅) = 5400

Factory overheads(6000+20℅) = 7200

Indirect materials and supplies= 96,200

Repairs and maintenance = 24,100

Plant service contracts= 37,000

Refurbishing cost= 89,100

Factory insurance = 8,200

Heat, light and power = 51,700

Indirect Labour = 120,000

Transportation = 25,000

Material= 80,000

Rent = 40,000

Security Cost = 15,000

Total of overheads= 8,06,100

C. If 60000 hours are used on production, overhead rate is:

Total sum of overhead/number of hours worked

= 754800/60000

= 12.58 $ per hour

= 754800/37770

= 19.98 $ per unit

Note: If we needs to calculate on units, then first find out how much units will produce in 60000 hours:

35000/55600x 60000

= 37770 units estimated in 60000 hours

Note: As overhead calculated ik Point No. B is estimated. And in part C, not mentioned that we have to calculate overhead rate for 60000 hours oj estimate Or actual cost, so we calculate overhead application rate for 60000 labour hours on actual overheads.

D. If ascertain time is 120 hours, then overhead is:

Actual labour hours = 55600

Ascertained hours = 120

Total overheads = 754800

Overheads for 120 hours for reaching the production of 35000 units = Actual hours/ expected hours x actual overheads

= 55600/120 x 754800  

= 34,97,24,000

Note: as time given is reduced to 463 times, so overhead cost will increase in directly proportionate to time reduction that is 463 times.